OpenStudy (anonymous):

find three consecutive positive integers such that the sum of the squares of the first and second equals the square of the third

OpenStudy (noelgreco):

Without doing the algebra, you could consider the smallest Pythagorean triple.

hartnn (hartnn):

let the smallest number be x so,other 2 numbers are x+1,x+2 so \(x^2+(x+1)^2=(x+2)^2\) solve for x.

OpenStudy (anonymous):

how?

hartnn (hartnn):

use \( (a+b)^2=a^2+2ab+b^2 \)

hartnn (hartnn):

u will get a quadratic in x.

OpenStudy (anonymous):

im sorry but can you help me step by step?

hartnn (hartnn):

\(x^2+(x+1)^2=(x+2)^2-->x^2+x^2+2x+1=x^2+4x+4\) ok?

hartnn (hartnn):

which gives \(x^2-2x-3=0\) which when u solve gives,x=-1,3 but as x is positive,x=3 only. so the numbers are 3,4,5.

OpenStudy (anonymous):

So next step is \[2x ^{2} +2x+1=x^{2}+4x+2\]

hartnn (hartnn):

u can subtract x^2+4x+4 from both sides.

OpenStudy (anonymous):

but how do i go from x^2-2x-3=0 to x=3?

hartnn (hartnn):

thats quadratic equation! u know any method to solve quadratic equations??

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

x^2+(x+1)^2=(x+2)^2 Use wolfram or something to solve for it lol.

hartnn (hartnn):

ok, x^2-2x-3=0 so, \(x^2-3x+x-3=0\) \(x(x-3)+1(x-3)=0\) \((x+1)(x-3)=0\) ok ? which step u did not understand?

OpenStudy (anonymous):

am9998, if you do not know how to solve quadratic equations, why are you working on this problem? Are you expected to use guess-and-check methods?

OpenStudy (anonymous):

No we were expected to know this. but my I didnt learn this.

OpenStudy (anonymous):

so far i understand but whats the next step?

hartnn (hartnn):

\((x+1)(x-3)=0\) so, x+1=0 OR x-3=0 so, x=-1 OR x=3 but x is positive so x=3 only.

hartnn (hartnn):

got it?

OpenStudy (anonymous):

yes thanks

hartnn (hartnn):

welcome :)